Jill borrows $17,000 from you today. She agrees to repay you in two equal amounts, the first to occur in 3 years from today and the other in 8 years from today. If the interest rate is 5.0% per annum compounding monthly, what will be the amount of each repayment?
Answer #1

Let 2 equal payments = x

Rate per month = 5%/12

PV of these payments = x/(1+5%/12)3*12 + x/(1+5%/12)8*12 = 17000

x* (0.860976+0.670877) = 17000

x = 11097.67

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well